OpenStudy (avanti):

Prove that sec(theta) / tan(theta) = csc(theta) please explain! note: there should be 3 steps

OpenStudy (psymon):

Do you know what sec(theta) and tan(theta) can be replaced with?

OpenStudy (avanti):

sec(theta) = 1/cos(theta) and tan(theta)= 1/cot(theta)

OpenStudy (psymon):

Right. But the one you replaced tan with isn't going to be quite as useful. Do you know what else is the same as tan(theta)?

OpenStudy (avanti):

sin(theta)/cos(theta) ?

OpenStudy (psymon):

There ya go. So now we would have: \[\frac{ 1 }{ \cos \theta } \div \frac{ \sin \theta }{ \cos \theta }\] Think you can see what will happen from here?

OpenStudy (avanti):

oh yeah! i got it thank you!

OpenStudy (psymon):

Awesome :3

OpenStudy (avanti):

do you think you can helo me with this one too? solve -sin^2 x= 2cosx - 2

OpenStudy (psymon):

Solve for x or prove?

OpenStudy (avanti):

solve please

OpenStudy (psymon):

Ah, alrighty. Lemme look.

OpenStudy (psymon):

Mkay, so the things to recognize is that we have a squared term and first power term. I see that and I think maybe its a quadratic. If it is, then we need to get everything to be sin or everything to be cos, as well as make sure they have the same angle. Well, they already have the same angle, so now we just need them to be the same trig function. Do you know how you might substitute sin^2(x) into something with cosines?

OpenStudy (avanti):

sin^2 x= (1-cos2x) / 2

OpenStudy (psymon):

Yes, but that gives us different angles. We want them all to just be x, not 2x. It's good you know that identity, though xD Well, we need to use this identity to help us get there: \[\sin ^{2}x + \cos ^{2}x = 1\] and then we solve for sin^2(x)

OpenStudy (avanti):

so sin^2x = 1- cos^2x ?

OpenStudy (psymon):

Yep, bingo. So this would give us: \[-(1-\cos ^{2}x) =2cosx - 2\] So now we have all we need to make this into a quadratic factoring equation, now we just need to set evertything equal to 0 and factor.

OpenStudy (avanti):

yeah that's the part i have trouble with.. i'm not quite sure how to make it into a quadratic function and then factor it

OpenStudy (psymon):

Did we at least set it up properly like a quadratic factoring problem?

OpenStudy (avanti):

No i think we have to have one side to = 0

OpenStudy (psymon):

No, I was just curious if you had doen that yourself yet or not xD

OpenStudy (psymon):

\[\cos ^{2}x - 2cosx +1 = 0\] Sorry bout that :P

OpenStudy (psymon):

Alright, so what I'm going to do is pretend cosx is simply just x. \[x ^{2} -2x +1 = 0\] Now we factor that.

OpenStudy (avanti):

so (x-1) (x-1)

OpenStudy (psymon):

Correct. So x = 1. Now we just re-replace x with cosx to get cosx = 1

OpenStudy (avanti):

so that means x =0. Is that my answer or should i leave like cosx=1 ?

OpenStudy (psymon):

No, x = 0 is the answer xD

OpenStudy (avanti):

Okay thank you so much for all your help, i really appreciate it! :)

OpenStudy (psymon):

Np :3